If 7^(2x)=3 What is the value of 7^(6x-1)?

Answer by lwsshak3(11628)

If 7^(2x)=3 What is the value of 7^(6x-1)?
***
2xlog7=log3
x=log3/2log7
..
7^(6x-1=7^(6log(3)/2log(7)-1)=1.5037...